I am stuck on what method I should use to solve this question, please help!!! A headline reads more than half of American safe federal taxes are too high the headline was based on a Random sample of 1026 adult Americans in which 534 stated the amount of federal taxes they had to pay was too high. B). Using .05 level of significance is there enough evidence to prove more than half of the adult Americans believe the amount of federal taxes they pays to high
Answer by Boreal(15235)

The point estimate is 0.5205
the se=sqrt (0.5*0.5/1026)=0.0156
the 95% CI is an interval 0.0156*1.96=0.0306
The CI is (0.4900, 0.5511)
Since 0.5 is in the confidence interval, there is insufficient evidence to reject Ho.
All the other parts of the hypothesis test can be filled in. The test is a one sample proportion, and reject for |z|>1.96, z's being the test statistic.
I did CI, but z=1.31 if it is done that way, and one still doesn't reject.